OpenStudy (anonymous):

Fan and Medal Find the roots of the polynomial equation. x^2 – x – 72 = 0 A. 9 and –8 B. –6 and 6 C. –2 and 16 D. –9 and 8

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

x^2 – x – 72 = 0 Factor the left hand side: x^2 - 9x + 8x - 72 = 0 Can you factor now?

OpenStudy (anonymous):

x+8(x-9) is what I got @ranga

OpenStudy (ranga):

Yes. (x+8)(x-9) = 0 Therefore, x + 8 = 0 or x - 9 = 0 x = -8 or x = 9. Those are the roots of the polynomial because those x values makes the polynomial zero.